The period of a particle in SHM is 8s. at t=0, it is at the mean position. the ratio of the distances trevelled by it in the 1st and 2nd second is ….....................

Grade:12

1 Answers

Arun
6 years ago

Dear student

 

Y¹= A sin w = A sin 2π\8 = A /√2 

Distance when t=2 sec is

Y²= A sin w= A sin π\2= A 

Distance covered in 2sec is A-A/√2 

Y¹/Y²= (A/√2) / (A-A/√2) 

= 1/√2-1

 

Regards

Arun (askIITians forum expert)